How to solve x+y=4, -2x+3y=7 by substitution
Answer by macston(5194)

x+y=4
x=4-y
-2x+3y=7 Substitute for x
-2(4-y)+3y=7
-8+2y+3y=7
-8+5y=7
5y=15
y=3 ANSWER 1:y=3
x+y=4
x+3=4
x=1 ANSWER 2:x=1
CHECK:
-2x+3y=7
-2(1)+3(3)=7
-2+9=7
7=7